# The Signals of the Doomsday

### v3

@@ 8.1 Cosmological inputs
$D_{\rm ly}=1.0\times 10^{8}\;\mathrm{ly},\quad D_{m}=D_{\rm ly}\times 9.4607\times 10^{15}\;\mathrm{m},$
@@ 9 Conclusions
The work presented here was motivated by an interplay between the LHC collider data and recent theoretical progress in the field theory in curved spacetime. The standard model precision data coming from the LHC indicate that we currently live in a false vacuum. While the LHC data indicate that the lifetime of the false vacuum is safely large, it has been shown that small primordial black holes can serve as the phase transition catalysers and significantly increase the transition probability. In fact, it might be possible that the bubbles of the true vacuum already exist in our visible universe. One of the signatures of the existence of such bubbles would be copious production of the heavy Higgs particles due to the vacuum mismatch inside and outside of the bubble. We first calculated the spectrum of Higgs particles produced in this way using the standard field theory techniques. We then used the HDECAY event generator to simulate their decay. We finally used Pythia to hadronize the decay products and obtain the photon and neutrino energy spectra as a long-range signature of the bubble production.
@@
The bubble wall’s velocity as seen by an external observer typically evolves in time as $v_{b}\sim t/\sqrt{t^{2}+R_{0}^{2}}$. Thus, for $t\gg R_{0}$, this velocity is almost the speed of light. However, it is never exactly equal to the speed of light. In fact, detailed calculations indicate that the bubble wall velocity might be limited to well below the speed of light [44]. In addition, the bubble wall can be viewed as a highly coherent state of a large number of the Higgs quanta. As such, it will strongly interact with virtually all the standard model particles. In the early universe, it will interact with the surrounding plasma, while at late times it will interact with other bubbles, stars, planets, interstellar and intergalactic gas, and finally even with the particles it creates itself. These interactions would further slow down the propagation of the wall. It is thus not very unreasonable to expect that these photons and neutrinos should be able to reach us before the bubble wall. No matter if this time delay is $10$ seconds or three years, if our gamma ray and neutrino observatories ever detect spectra corresponding to Figures 6 and 7, they might in fact represent the signals of the doomsday.
Since the particle production mechanism we relied on here resembles the Unruh effect, which in turn requires a non-zero proper acceleration of the wall, we included only particles produced until the terminal velocity is reached. After that, some other mechanisms might keep producing particles, for example because huge energy is dumped into the environment due to friction.
@@
In broader context, a late time first‐order electroweak transition is expected to leave a multi‐messenger imprint— possibly a stochastic gravitational‐wave background, subtle distortions in the CMB and 21-cm line, and potential signals in gamma-ray, cosmic-ray and pulsar-timing data—that upcoming facilities are well poised to detect. These direct and indirect channels provide the most robust avenues for detecting the bubbles of the true vacuum. In a follow-up study, we will explore the detectability and present detailed forecasts for both current and next-generation observatories.
@@
We note that the crucial assumption here is that there is no new physics beyond the standard model. New physics could potentially modify the Higgs potential and render our vacuum stable. Another possibility is that small primordial black holes do not exist. Since they are necessary catalysers, their absence in an appropriate mass range [18] would perhaps invalidate this doomsday scenario.
@@ Acknowledgments
10 Appendix-A
@@
$t=R_{0}\sinh\left(\frac{\tau}{R_{0}}\right),\quad r=R_{0}\cosh\left(\frac{\tau}{R_{0}}\right).$
@@
$R_{0}^{2}\cosh^{2}\left(\frac{\tau}{R_{0}}\right)-R_{0}^{2}\sinh^{2}\left(\frac{\tau}{R_{0}}\right)=R_{0}^{2}.$
@@
$u^{\mu}=\left(\frac{dt}{d\tau},\frac{dr}{d\tau}\right)=\left(\cosh\left(\frac{\tau}{R_{0}}\right),\sinh\left(\frac{\tau}{R_{0}}\right)\right).$
@@
$u^{\mu}u_{\mu}=\left(\frac{dt}{d\tau}\right)^{2}-\left(\frac{dr}{d\tau}\right)^{2}=\cosh^{2}\left(\frac{\tau}{R_{0}}\right)-\sinh^{2}\left(\frac{\tau}{R_{0}}\right)=1,$
@@
$a^{\mu}=\frac{du^{\mu}}{d\tau}=\left(\frac{1}{R_{0}}\sinh\left(\frac{\tau}{R_{0}}\right),\frac{1}{R_{0}}\cosh\left(\frac{\tau}{R_{0}}\right)\right).$
@@
$a^{\mu}a_{\mu}=\left(\frac{1}{R_{0}}\sinh\left(\frac{\tau}{R_{0}}\right)\right)^{2}-\left(\frac{1}{R_{0}}\cosh\left(\frac{\tau}{R_{0}}\right)\right)^{2}=-\frac{1}{R_{0}^{2}}.$
@@
11 Appendix-B
Validation of the Thin-Wall Approximation
  • The true Higgs background is a smooth kink of width $\ell_{\rm wall}\sim 1/\mu$, interpolating between $\varphi_{\rm false}$ and $\varphi_{\rm true}$ over $\xi=(r-R(t))/\ell_{\rm wall}\in[-\tfrac{1}{2},\tfrac{1}{2}]$. In the thin-wall limit one sends $\ell_{\rm wall}\to 0$ at fixed surface tension $\sigma$, or equivalently requires
    $\varepsilon\;=\;\frac{\ell_{\rm wall}}{R(t)}\;\ll\;1.$
    For our fiducial nucleation parameters, $\ell_{\rm wall}\approx 1/\mu\approx 1.4\times 10^{3}\,M_{\rm Pl}^{-1}$ and $R_{0}=2\times 10^{4}\,M_{\rm Pl}^{-1}$, giving
    $\varepsilon=\frac{1/\mu}{R_{0}}=\frac{1.4\times 10^{3}}{2\times 10^{4}}\approx 0.07\;\ll\;1,$
    so curvature effects across the wall thickness are negligible.
  • In this regime the smooth mass profile $m^{2}(r,t)=\mu^{2}\,f(\xi)$ may be replaced by a sharp Heaviside step:
    $m^{2}(r,t)\approx\begin{cases}0\,,&r<R(t),\\ \mu^{2}\,,&r>R(t).\end{cases}$
    Finite-thickness corrections enter only at $\mathcal{O}(\varepsilon)$, modifying the Bogoliubov coefficients by $\mathcal{O}(\ell_{\rm wall}\,\mu)\sim\mathcal{O}(1)$ prefactors, but they cannot turn the exponentially suppressed factor $\exp(-4\mu/a)$ into an $\mathcal{O}(1)$ effect when $a<\mu$.
  • An independent WKB/adiabaticity check further confirms this. We define
    $\epsilon\;=\;\frac{1}{\omega^{3}}\Bigl{|}\tfrac{d}{dt}\,m^{2}\Bigr{|},\quad\omega\sim\mu.$
    By the chain rule,
    $\frac{d}{dt}m^{2}\sim m\,\dot{R}\,\frac{dm}{d\xi}\,\frac{1}{\ell_{\rm wall}}\sim\mu\times 1\times 1\times\mu=\mu^{2},$
    and thus
    $\epsilon\sim\frac{\mu^{2}}{\mu^{3}}=\frac{1}{\mu}\;\ll\;1.$
    Since $\dot{R}\lesssim 1$, $m\approx\mu$, $\tfrac{dm}{d\xi}\sim\mathcal{O}(1)$ (in units of wall widths), and $\ell_{\rm wall}\sim 1/\mu$, whenever $\mu\gg a$ we have $\epsilon\ll 1$, so mode mixing and particle production remain exponentially suppressed, $\sim e^{-\mu/a}$.
  • We conclude that using a thin wall approximation in our context is a controlled approximation at the few-percent level ($\varepsilon\approx 0.07$), and that finite-thickness effects cannot restore significant continuous production when $a<\mu$.

### v4

@@ 8.1 Cosmological inputs
$D_{\rm ly}=1.0\times 10^{8}\;\mathrm{ly},\quad D_{m}=D_{\rm ly}\times 9.4607\times 10^{15}\;\mathrm{m},$ (8.6)
@@ 9 Conclusions
The work presented here was motivated by an interplay between the LHC collider data and recent theoretical progress in the field theory in curved spacetime. The standard model precision data coming from the LHC indicate that we currently live in a false vacuum. While the LHC data indicate that the lifetime of the false vacuum is safely large, it has been shown that small primordial black holes can serve as the phase transition catalysers and significantly increase the transition probability. In fact, it might be possible that the bubbles of the true vacuum already exist in our visible universe. One of the signatures of the existence of such bubbles would be copious production of the heavy Higgs particles due to the vacuum mismatch inside and outside of the bubble. We used the standard field theory techniques to calculate the spectrum of Higgs particles produced this way. Ideally, one would want to use the event generators to simulate the Higgs decay and subsequent evolution of the decay products. However, with the current state-of-the-art in the field, this is not possible. The mass of the Higgs particle in the new vacuum is very large ($\sim 10^{15}\,\mathrm{GeV}).$ At these scales, the perturbation theory certainly breaks down, and the results from the standard event generators cannot be extrapolated. Moreover, there are indications that far above the electroweak breaking scale the Higgs does not even behave as a well defined particle, but rather as a resonance [38]. Fortunately, we can reasonably argue that most of the energy invested into the Higgs particle production will eventually end up in photons and neutrinos. The reason is that the Higgs is a neutral particle and will decay into an equal number of particles and antiparticles. Some of them will be unstable and will further decay, and some of them will hadronize and then decay. As the final product we will get mostly photons, neutrinos, leptons and anti-leptons. Some of these leptons and anti-leptons will annihilate giving again photons. Thus, most of the energy will go into a long-range signature - photons and neutrinos.
@@
The bubble wall’s velocity as seen by an external observer typically evolves in time as $v_{b}\sim t/\sqrt{t^{2}+R_{0}^{2}}$. Thus, for $t\gg R_{0}$, this velocity is almost the speed of light. If this remains true at all times, a very large number of particles will be produced due to the vacuum mismatch since the proper acceleration never ceases, but we will have no chance to observe them before the wall hits us. However, the wall velocity is never exactly equal to the speed of light. In fact, detailed calculations indicate that the bubble wall velocity might be limited to well below the speed of light [20]. In addition, the bubble wall can be viewed as a highly coherent state of a large number of the Higgs quanta. As such, it will strongly interact with virtually all the standard model particles. In the early universe, it will interact with the surrounding plasma, while at late times it will interact with other bubbles, stars, planets, interstellar and intergalactic gas, and finally even with the particles it creates itself. These interactions would further slow down the propagation of the wall. It is thus not very unreasonable to expect that these photons and neutrinos should be able to reach us before the bubble wall.
If the friction with the environment exists, then the proper acceleration will cease at some point. Since the particle production due to vacuum-mismatch mechanism is analogous to an Unruh-type emission that operates only while the bubble wall is accelerating, our calculated spectrum accounts for particle production only up to the point where the wall attains its terminal velocity. After acceleration ceases, the mismatch channel shuts off. Nevertheless, the wall continues to transfer enormous energy into the environment through friction during the terminal phase, and this energy is rapidly thermalized in the shocked layer behind the wall. As we have shown, this thermalization process itself becomes an efficient source of additional particles, complementing the vacuum-mismatch production even after the wall has stopped accelerating. This process produces an enormous amount of particles in a very short period of time, which means that the period of friction does not have to last for very long. For our conservative benchmarks the total neutrino yield is some orders of magnitude smaller than that of a typical core–collapse supernova, but the characteristic energies are much higher, and a bubble nucleated very close to us would still appear as an exceptionally bright, short–duration burst.
If friction continues for some extended period of time, a huge vacuum energy density which is converted into thermal radiation may exceed the Planckian energy density in a thin layer around the bubble. In that case, the resulting spacetime geometry may likely resemble a (spherical) black membrane. However, such extended objects are unstable and will most likely decay into a large number of smaller black holes. If that indeed happens, then the phase transition might effectively be quenched, and the signals might reach us without the subsequent catastrophe.
@@
In a broader context, a late time first‐order electroweak transition is expected to leave a multi‐messenger imprint— possibly a stochastic gravitational‐wave background, subtle distortions in the CMB and 21-cm line, and potential signals in gamma-ray, cosmic-ray and pulsar-timing data—that upcoming facilities are well poised to detect. These direct and indirect channels provide the most robust avenues for detecting the bubbles of the true vacuum. In a follow-up study, we will explore the detectability and present detailed forecasts for both current and next-generation observatories.
@@
We note that the crucial assumption here is that there is no new physics beyond the standard model. New physics could potentially modify the Higgs potential and render our vacuum stable. Another possibility is that small primordial black holes do not exist. Since they are necessary catalysers, their absence in an appropriate mass range [36] would perhaps invalidate this doomsday scenario. We also stress that even if Higgs-instability regions are seeded in the Universe, their fate need not be “doomsday” in all circumstances: De Luca–Kehagias–Riotto [41] argue that in the thin-wall regime such dangerous patches can collapse due to Higgs backreaction (often leaving black holes), rather than expanding and eating the Universe, which can substantially relax the corresponding inflationary Hubble bound.
@@ Acknowledgments
Appendix A Appendix A: Proper Acceleration of the Bubble Wall
@@
$t=R_{0}\sinh\left(\frac{\tau}{R_{0}}\right),\quad r=R_{0}\cosh\left(\frac{\tau}{R_{0}}\right).$ (A.1)
@@
$R_{0}^{2}\cosh^{2}\left(\frac{\tau}{R_{0}}\right)-R_{0}^{2}\sinh^{2}\left(\frac{\tau}{R_{0}}\right)=R_{0}^{2}.$ (A.2)
@@
$u^{\mu}=\left(\frac{dt}{d\tau},\frac{dr}{d\tau}\right)=\left(\cosh\left(\frac{\tau}{R_{0}}\right),\sinh\left(\frac{\tau}{R_{0}}\right)\right).$ (A.3)
@@
$u^{\mu}u_{\mu}=\left(\frac{dt}{d\tau}\right)^{2}-\left(\frac{dr}{d\tau}\right)^{2}=\cosh^{2}\left(\frac{\tau}{R_{0}}\right)-\sinh^{2}\left(\frac{\tau}{R_{0}}\right)=1,$ (A.4)
@@
$a^{\mu}=\frac{du^{\mu}}{d\tau}=\left(\frac{1}{R_{0}}\sinh\left(\frac{\tau}{R_{0}}\right),\frac{1}{R_{0}}\cosh\left(\frac{\tau}{R_{0}}\right)\right).$ (A.5)
@@
$a^{\mu}a_{\mu}=\left(\frac{1}{R_{0}}\sinh\left(\frac{\tau}{R_{0}}\right)\right)^{2}-\left(\frac{1}{R_{0}}\cosh\left(\frac{\tau}{R_{0}}\right)\right)^{2}=-\frac{1}{R_{0}^{2}}.$ (A.6)
@@
Appendix B Appendix B: Numerical Procedure for Particle Production due to Vacuum Mismatch
To get the total number of particles from (6.8) we discretize proper time $\tau$ and integrate the coupled system $\{y(\tau),t(\tau),R(\tau),N_{k=0}^{\rm(int)}\}$ defined by (6.3)–(6.8) using a classical fourth-order Runge–Kutta (RK4) method with adaptive steps. The state is initialized at nucleation by $y(0)=0$ $(v=0)$, $t(0)=0$, $R(0)=R_{0}$, and $N^{(\text{int})}_{k=0}(0)=0$. At each substep we evaluate the raw proper acceleration
$\alpha_{\rm raw}(\tau)=A-\frac{2}{R(\tau)}-B\,\sinh y(\tau).$ (B.1)
$\frac{dy}{d\tau}=\alpha_{\rm raw},\qquad\frac{dt}{d\tau}=\cosh y,\qquad\frac{dR}{d\tau}=\sinh y,$ (B.2)
and accumulate the produced quanta using
$\frac{dN_{k=0}}{d\tau}=N_{k=0}(\tau)\,4\pi\,R(\tau)^{2}\,\sinh y(\tau),\qquad N_{k=0}(\tau)=\left[\frac{(\omega_{+}+\omega_{-})^{2}}{(\omega_{+}-\omega_{-})^{2}}e^{\frac{4\omega_{+}}{\alpha_{raw}}}-1,\right]^{-1}.$ (B.3)
$N_{k=0}(\tau)\;\simeq\;\left[\frac{(\omega_{+}+\omega_{-})^{2}}{(\omega_{+}-\omega_{-})^{2}}\right]^{-1}e^{-4\mu/\alpha_{\rm raw}}\quad\text{when}\quad\frac{4\mu}{\alpha_{\rm raw}}\gtrsim 200,$ (B.4)
$d\tau\;\sim\;\frac{0.05}{\max\!\big(|\alpha_{\rm raw}|,10^{-6}\big)},$ (B.5)
Stopping criterion and diagnostics. The integration stops when the proper time first satisfies $\tau=\tau_{\rm term}$ for the scenario. For accuracy control, we repeat each run with the initial $d\tau$ halved and demand relative changes $\lesssim 10^{-4}$ in $\{\tau_{\rm term},R_{\rm fin},N_{k=0}^{\rm(int)}\}$. As additional consistency checks we verify: (a) In the planar, late–time regime where $R\to\infty$ (so $2\sinh y/R^{2}\to 0$), Eq. (5.31) reduces to $d\alpha/d\tau\simeq-B\cosh y\,\alpha\leq-B\alpha$. Therefore $\alpha(\tau)\leq\alpha(\tau_{0})\,e^{-B(\tau-\tau_{0})}$ holds for $\tau\geq\tau_{0}$ once the curvature term becomes negligible. We verify numerically that the inequality is satisfied after the time when $2\sinh y/R^{2}\ll B\cosh y\,\alpha$; (b) monotonicity of $R(\tau)$; and (c) that $N_{k=0}(\tau)$ vanishes whenever $\alpha_{\rm raw}\leq 0$. To illustrate, we consider three ultra-relativistic terminal deficits $\delta\equiv 1-v_{\rm term}$ and use the given friction pairs $(\eta,\tau_{\rm term})$ for each scenario; this fixes $B=\eta/\sigma$ and thus the proper-time dynamics via (6.4)–(6.8). At $\tau=\tau_{\rm term}$ we report: the final proper time $\tau_{\rm fin}$, radius $R_{\rm fin}$, and the integrated yield $N_{\rm tot}^{\rm(int)}$.
$\gamma_{\rm term}\,v_{\rm term}\;=\;\frac{\Delta V}{\eta},$ (B.6)
$\eta(\delta)\;=\;\frac{\Delta V}{\gamma_{\rm term}\,v_{\rm term}}.$ (B.7)
Appendix C Appendix C: Justification of the cutoffs for vacuum–mismatch and thermal production
In this appendix we justify cutting off the vacuum–mismatch contribution at $\tau_{\rm term}$, while evolving the thermal sector up to $\tau_{\rm final}=5\tau_{\rm term}$. We use the Higgs benchmark and notation of Sec. 6, with parameters in Eq. (6.1) and proper acceleration $\alpha(\tau)$ given by Eq. (6.4).
Vacuum–mismatch production: exponential shutoff
The vacuum–mismatch occupancy $N_{k=0}(\tau)$ is given in Eq. (6.6), and the integrated zero–mode yield $N_{k=0}^{(\text{int})}(\tau)$ in Eq. (6.8). For small positive proper acceleration $\alpha(\tau)$ one may use the asymptotic form
$N_{k=0}(\tau)\simeq\left[\frac{(\omega_{+}+\omega_{-})^{2}}{(\omega_{+}-\omega_{-})^{2}}\right]^{-1}\exp\!\left(-\frac{4\omega_{+}}{\alpha(\tau)}\right),\qquad\omega_{+}=\sqrt{\mu^{2}+k^{2}},\,\omega_{-}=\sqrt{M^{2}+k^{2}},$ (C.1)
so that the production rate is exponentially suppressed once $\alpha(\tau)$ becomes small.
For the Higgs benchmark,
$\mu=7.16\times 10^{-4}M_{\rm P},\qquad R_{0}=2\times 10^{4}M_{\rm P}^{-1},$
so at nucleation $\alpha(0)\simeq 1/R_{0}=5.0\times 10^{-5}M_{\rm P}$. In the late–time planar regime the proper acceleration decays on the timescale $\tau_{\rm term}$, and it is convenient to model it as
$\alpha(\tau)\simeq\alpha(0)\,e^{-\tau/\tau_{\rm term}}.$ (C.2)
At $\tau=3\tau_{\rm term}$ this gives
$\alpha(3\tau_{\rm term})=\alpha(0)e^{-3}\simeq 2.49\times 10^{-6}M_{\rm P},$ (C.3)
so that
$x(3\tau_{\rm term})\equiv\frac{4\mu}{\alpha(3\tau_{\rm term})}=\frac{4\times 7.16\times 10^{-4}}{2.49\times 10^{-6}}\simeq 1.15\times 10^{3},$ (C.4)
and
$N_{k=0}(3\tau_{\rm term})\sim\exp[-x(3\tau_{\rm term})]=\exp(-1.15\times 10^{3})\simeq 10^{-4.997\times 10^{2}}\simeq 2.2\times 10^{-500}.$ (C.5)
Already at $3\tau_{\rm term}$ the $k=0$ occupancy is effectively zero.
The late–time contribution from the vacuum–mismatch channel between $\tau_{\rm term}$ and $3\tau_{\rm term}$ is
$\Delta N_{k=0}^{(\text{int})}(3\tau_{\rm term})=\int_{\tau_{\rm term}}^{3\tau_{\rm term}}\!{\rm d}\tau\,N_{k=0}(\tau)\;4\pi R^{2}(\tau)\,\sinh y(\tau).$ (C.6)
For the most ultra–relativistic benchmark ($\delta=10^{-10}$ in Table 1), the numerical evolution up to $\tau_{\rm term}=\sigma/\eta$ gives
$R(\tau_{\rm term})\simeq 7.1\times 10^{8}\,M_{\rm P}^{-1},\qquad\gamma(\tau_{\rm term})\simeq 7.1\times 10^{4}.$ (C.7)
During the subsequent interval the wall remains ultra–relativistic and expands at $v\simeq 1$, so a conservative bound is
$R^{2}(\tau)\;\lesssim\;10^{19},\qquad\sinh y(\tau)\;\lesssim\;\gamma_{\rm term}\;\sim\;7\times 10^{4},\qquad(\tau_{\rm term}\leq\tau\leq 3\tau_{\rm term}),$ (C.8)
which implies
$4\pi R^{2}(\tau)\,\sinh y(\tau)\;\lesssim\;4\pi\times 10^{19}\times 7\times 10^{4}\sim 10^{25}.$ (C.9)
Over the same interval we may bound $N_{k=0}(\tau)$ from above by its value at $\tau=3\tau_{\rm term}$, since $\alpha(\tau)$ continues to decrease. Using $N_{k=0}(3\tau_{\rm term})\simeq 2.2\times 10^{-500}$ we obtain
[…]
The width of the interval is
$\Delta\tau_{(3)}=3\tau_{\rm term}-\tau_{\rm term}=2\tau_{\rm term}\simeq 1.4\times 10^{9}\,M_{\rm P}^{-1},$ (C.11)
so that
$\Delta N_{k=0}^{(\text{int})}(3\tau_{\rm term})\lesssim 10^{-475}\times 10^{9}\sim 10^{-466}.$ (C.12)
This should be compared with the integrated zero–mode yields obtained in Sec. 6 and summarised in Table 1,
$N_{k=0}^{(\text{int})}(\delta=10^{-8})\simeq 1.05\times 10^{8},\quad N_{k=0}^{(\text{int})}(\delta=10^{-9})\simeq 3.53\times 10^{9},\quad N_{k=0}^{(\text{int})}(\delta=10^{-10})\simeq 1.14\times 10^{11}.$ (C.13)
The late–time contribution from $\tau\in[\tau_{\rm term},3\tau_{\rm term}]$ is therefore suppressed by roughly $10^{474}$$10^{477}$ compared to the values in Table 1, and is numerically irrelevant. Since $\alpha(\tau)$ continues to decay on the timescale $\tau_{\rm term}$ for $\tau>3\tau_{\rm term}$, extending the upper limit further (e.g. to $5\tau_{\rm term}$) only decreases $N_{k=0}(\tau)$ by additional thousands of orders of magnitude, and the correction to $N_{k=0}^{(\text{int})}$ remains negligible. This justifies cutting off the vacuum–mismatch calculation at $\tau_{\rm term}$: the contribution is dominant for $\tau\lesssim\tau_{\rm term}$, while for $\tau\gtrsim\tau_{\rm term}$ the Unruh–like channel governed by $\alpha(\tau)$ is exponentially suppressed.
Thermal production: sensitivity to the integration time
For the thermal channel we use the same wall trajectories $R(\tau),y(\tau)$ but count the number of quanta in the heated shell, assuming an approximately thermal distribution. The instantaneous rate (cf. Sec. 7) has the schematic form
[…]
with temperature $T(\tau)\propto\rho_{\rm th}(\tau)^{1/4}$ determined by the thermal energy density $\rho_{\rm th}(\tau)$. The key difference from the vacuum–mismatch case is that this source does not require a nonzero acceleration: as long as the shell remains hot ($T(\tau)>0$), and the wall continues to sweep volume, thermal production continues.
These features are easily illustrated in a simple late–time toy model. Assume that the wall has already reached a fixed terminal rapidity $y_{\rm term}$, so that $R(\tau)\simeq v_{\rm term}\tau$ and $\sinh y(\tau)\simeq\sinh y_{\rm term}$, and that the temperature is approximately constant over the interval of interest. Then
[…]
Normalising $N_{\rm th}(\tau_{\rm term})$ to unity in this toy model, extending the upper limit from $\tau_{\rm term}$ to $3\tau_{\rm term}$ and $5\tau_{\rm term}$ gives
$\frac{N_{\rm th}(3\tau_{\rm term})}{N_{\rm th}(\tau_{\rm term})}\simeq 3^{3}=27,\qquad\frac{N_{\rm th}(5\tau_{\rm term})}{N_{\rm th}(\tau_{\rm term})}\simeq 5^{3}=125.$ (C.16)
Thus, even with a strictly constant temperature one naturally expects enhancements of order $\mathcal{O}(10)$ when integrating to $3\tau_{\rm term}$ and of order $\mathcal{O}(10^{2})$ when integrating to $5\tau_{\rm term}$.
In the full energy–deficit setup of the main text, the thermal energy density $\rho_{\rm th}(\tau)$, and hence $T(\tau)$, continue to grow for some time beyond $\tau_{\rm term}$ due to the ongoing energy deficit between the frictionless and friction–limited walls. This further amplifies the sensitivity of $N_{\rm th}$ to the upper integration limit, and in practice we find that evolving to $\tau_{\rm final}=5\tau_{\rm term}$ can increase the final thermal multiplicities by several orders of magnitude compared to integrating only up to $\tau_{\rm term}$.